Kingbird Corporation’s charter authorized issuance of 96,000 shares of $10 par value common stock and 47,700 shares of $50 preferred stock. The following transactions involving the issuance of shares of stock were completed. Each transaction is independent of the others.

1. Issued 357 shares of common and 95 shares of preferred for a lump sum amounting to $9,700. The common had been selling at $15 and the preferred at $61.

2. Issued 210 shares of common and 47 shares of preferred for equipment. The common had a fair value of $17 per share; the equipment has a fair value of $7,000.

Record the transactions listed above in journal entry form
Answer #1

Solution:

Fair value of common share issued = 357*15 = $5,355

Fair value of preferred stock issued = 95*$61 = $5,795

Lumpsum issue price = $9,700

Amount allocated to common stock = $9,700 * $5,355 / ($5,355 + $5,795) = $4,659

Amount allocated to common stock = $9,700 * $5,795 / ($5,355 + $5,795) = $5,041

Kingbird Corporation
Journal Entries
Event Particulars Debit Credit
1 Cash Dr $9,700.00
            To Common stock $3,570.00
            To Paid in capital in excess of par - common stock $1,089.00
            To Preferred stock $4,750.00
            To Paid in capital in excess of par - Preferred stock $291.00
(To record issue of shares)
2 Equipment Dr $7,000.00
            To Common stock $2,100.00
            To Paid in capital in excess of par - common stock $1,470.00
            To Preferred stock $2,350.00
            To Paid in capital in excess of par - Preferred stock $1,080.00
(To record purchase of equipment by issue of shares)